MODS1 AGw Commissioning at LDG: 28 September 2010 UTC

Observers: R. Pogge (OSU), R. Stoll (OSU)

Telescope Operator: DHuerta (LBTO)

AO Support: D. Miller (LBTO)

Telescope Support: M. Midkiff (LBTO)

Instrument Support: J. Morris (LBTO)

Summary

Heavy high clouds moving NW at sunset, maybe clearing behind them, but looks to be lots of sucker holes ahead. Lost almost 2h at the start to a telescope azimuth encoder problem.

Doug Miller was up for AO work during the day, but he stayed up late with us to work on getting the AGw adaptive optics tuned up. Doug also got his scripts running for automating data acquisition for measuring the stage transformations, including debugging interactions with MODS.

Mostly overcast all night, very little on-sky time for our last night. Did some internal calibration work in a dark dome, but otherwise we quit at 10:40 UTC as clouds were still pretty bad, and we need to get up to button up the instrument tomorrow before leaving.

Preparations

Instrument Mounting Bolt Inspection

Mike Midkiff inspected the MODS mounting bolts at horizon pointing, checking the bottom-center bolt at each 60-degree rotation interval.

All six bolts were again tight - none were loose.

Details

Dewar Fill

Filled at 18:30MST, still showing at least 24h hold time on both CCDs.

This is the last filling for the time being.

Opening Up

02:20 - Clouds thinning, opening up. RH=50% T=12C WS=6m/s from the E with gusts to 14 m/s

Still a little murky out high up, and wind is out of the east with a few gusts. Cooler than last night at this time.

02:48 - Pointing Corrections: IE=-6, CA=15

02:54 - AZ drive fault, DGH clearing...

02:55 - Faint star ACT0253F, guide preset

02:58 - AZ drive fault again... Mike M called in to help at 03:02.

03:11 - getting 100% torque on AZ rotation. DGH and MM working on it...

03:47 - Sending preset. Rotator was turned off

Can hear the speed wind up and down as the AZ drive slews around, then we see red on the MCS control panel.

Fault again, David calling for the cavalry (Dave Ashby).

03:56 - David closing the shutter doors during the diagnosis because of clouds coming back...

04:35 - Azimuth encoder problem, problem encoder disabled, running on the others.

Clouds are back, moon is rising. Nothing to look at for now. We're still closed.

05:00 - Still clouds, not yet open. Will check on the half hour (sooner if miraculous clearing)

Calibrations

05:45 - still cloudy (worse than before).

Taking some calibration data - flat fields, biases, etc. while the dome is dark.

Red Spectral Flats: Red Grating CenLam=7750, LS5x60x0.6 slit, vflat=5.0, exptime=45sec, 10x (mods1r.20100928.0002-12)

Ne Lamp: Same setting, exptime 1s + ND1.5 filter, 3x (mods1r.20100928.0014-0016) Ar Lamp: Same, 1s and 2s, only a couple of lines saturated (mods1r.20100928.0017-0019)

06:25 - Doug Miller testing scripts that use the "doMODS" engineering command line to take MODS data from engineering scripts (IRAF, IDL, etc.). Mostly blue channel so he doesn't mess with our current red settings.

Still cloudy, maybe a sucker hole on its way

07:00 - Finally getting a break in the clouds - RH=40% T=11C Dewpoint=-2.6C WS=5-8m/s, gusts to 12m/s

07:05 - Doors open, ready to start over again.

07:06 - acquire preset for a bright pointing star - WT10-259

07:12 - Revised pointing corrections: CA=+9 IE=-22

Note, previously the pointing model had not been loaded.

07:15 - Faint pointing/guide setup star: ACT-0293F, guide preset

Doug running the IDL program to try to tweak the collimation. Acquisition succeeded

07:18 - WFS working in IDL, showing improvement. Seeing estimates 0.8-1.0 arcsec. DIMM reports worse.

Coordinate Transformations

07:25 - Guide Preset to Stone-A2

Guide acquire failed in bright background - Doug witnessed. Clearly some kind of sky threshold fault. We wonder what the algorithm is and why it is so sensitive.
See Issue 2893.

finally got an acquire with the ND1.0 filter in the guider, but that is no guarantee...

Acquiring science camera images to estimate good exposure times to avoid saturation in this seeing.

We got one image (mods1r.20100928.0022.fits) of Stone-A2 in SDSS r in 30s integration and 0.6-arcsec FWHM images. Background, however, is high, and this particular image has the guide probe encroaching upon the science field (by design).

Rats! Clouds coming back (sucker hole) - guide star fade and loss of lock while setting up to take a second image of the field with a different guide star preset (test of the transformation sequence by-hand).

07:55 - Clouds back with a vengeance. David now closing.

Wind jumped to sustained 9m/s, gusts to 12m/s, temp still steady at 11C, humidty around 46%. Broken flocculent high clouds illuminated by bright moonlight.

Doug is using the time to get the scripts for the transformations working, taking the opportunity of having a live MODS available to debug the instrument communications modules.

08:32 - High clouds cover nearly the entire sky, and are obscuring the moon.

10:00 - Still cloudy and closed up. Getting pretty hopeless for the rest of the night given all the restart we'd need to do and when twilight begins (such is the problem of a visible-light instrument on an 8.4m telescope).

We're taking 20min dark exposures inside the instrument with the chamber lights off and the instrument closed up (dark hatch, dark mask, and calibration tower in the beam). This is the first time we'll have taken "darks" in the same chamber environment as observers may use, and which we cannot reproduce in the lab. It'll also show us what LEDs and other light leaks are still on inside the instrument.

10:41 - Not getting any better, so we're calling it a night.

-- RichardPogge - 28 Sep 2010
Topic revision: r4 - 29 Sep 2010, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback